Let A and B be two events associated with an experiment and suppose that P(A)=0.4 while P(AUB)=0.7. Let


P(B)=P


a) For what choice of P are A and B mutually exclusive?


b) For what choice of P are A and B independent?

1
Expert's answer
2022-04-20T02:26:38-0400

a) For mutually exclusive events the following equality is true:

"P(A\cup B)=P(A)+P(B),\\
P=P(B)=P(A\cup B)-P(A)=0.7-0.4=0.3."


b) For joint events

"P(A\cup B) =P(A)+P(B)-P(A\cap B)."

For independent events

"P(A\cap B)=P(A)\cdot P(B)."

So,

"P(A\cup B) =P(A)+P(B)-P(A)\cdot P(B),\\
0.7=0.4+P-0.4P,\\
0.6P=0.3,\\
P=0.5."